# Learning What Matters: Prioritized Concept Learning via Relative Error-driven Sample Selection

### v1

@@ 4.1 Experimental Setup
Datasets and Models. To demonstrate effectiveness and generalizability of our approach across different scales of instruction-tuning (IT) data, we conduct experiments on two IT datasets: Visual-Flan-191K (Xu et al., 2024) and the larger-scale LLaVA-665K (Liu et al., 2023b) containing $\sim 0.6$ million samples. For the target VLMs, we use LLaVA-v1.5-7B (Liu et al., 2023b), following prior works. Additionally, we use LLaVA-v1.5-13B (Liu et al., 2023b) to test scalability and Qwen2-VL-7B (Wang et al., 2024) to test architecture generalization towards newer VLMs.
Implementation Details. Following standard protocol(Lee et al., 2024), we adopt LoRA (Hu et al., 2021) for training using the official hyperparameters from LLaVA-1.5. For the accuracy-based variant, we estimate cluster-wise accuracy using an LLM judge that compares the VLM output to ground-truth answers for samples in given cluster—though this is not required for our loss-based variant. Our setup follows standard evaluation protocols from prior work, ensuring consistency and fair comparison. Further details are provided in Appendix A.
Baselines. We compare PROGRESS against strong baselines spanning five major categories: (1) scoring function based methods (CLIP-Score, EL2N(Paul et al., 2021), Perplexity (Marion et al., 2023)); (2) deduplication-based selection (SemDeDup (Abbas et al., 2023)); (3) graph-based methods (D2-Pruning (Maharana et al., 2024)); (4) learned static selectors (Self-Filter (Chen et al., 2024a)); and (5) concept-diversity approaches (COINCIDE (Lee et al., 2024), Self-Sup (Sorscher et al., 2022)). We also include Random—a simple yet competitive baseline shown to perform well due to its diversity—and Full-Finetune, representing the performance upper bound with full supervision. Details for baselines follow previous work (Lee et al., 2024) and are provided in Appendix A.
Evaluation Benchmark. We evaluate our approach on a diverse suite of 14 vision-language benchmarks targeting different skills: perceptual reasoning (VQAv2 (Goyal et al., 2017), VizWiz (Gurari et al., 2018)), textual reasoning (TextVQA (Singh et al., 2019)), compositional reasoning (GQA (Hudson & Manning, 2019)), object hallucinations (POPE (Li et al., 2023b)), multilingual understanding (MMBench-cn (Liu et al., 2024a), CMMMU (Ge et al., 2024)), instruction-following (LLaVA-Bench(Liu et al., 2023b)), fine-grained skills (MME (Liang et al., 2024), MMBench-en (Liu et al., 2024a), SEED (Li et al., 2023a)), and scientific questions and diagrams (SQA-I (Lu et al., 2022), AI2D (Kembhavi et al., 2016), ChartQA (Masry et al., 2022)). See Appendix A for details.
@@
Table 1: Comparison of coreset selection techniques for training LLaVA-v1.5 on the LLaVA-665K dataset using 20% sampling ratio.Methods highlighted in orange require additional reference VLMs and 100% dataset annotations for coreset selection, while methods highlighted in light green do not require either. The benchmark results are highlighted with best and second best models within the respective categories (i.e, with and without utilizing additional information). The best and the second best relative score are in bold and underlined, respectively.
Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA- SEED AI2D ChartQA CMMMU Rel. (%)
en cn Wild
LLaVA-v1.5-7B
0   Full-Finetune 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 67.0 56.4 16.4 22.1 100
1   Self-Sup 74.9 59.5 46.0 67.8 49.3 83.5 1335.9 61.4 53.8 63.3 62.5 52.9 16.1 23.4 94.6
1   Self-Filter 73.7 58.3 53.2 61.4 52.9 83.8 1306.2 48.8 45.3 64.9 60.5 48.7 14.1 19.8 90.1
3   EL2N 76.2 58.7 43.7 65.5 53.0 84.3 1439.5 53.2 47.4 64.9 61.8 49.3 16.5 23.9 93.4
4   SemDeDup 74.2 54.5 46.9 65.8 55.5 84.7 1376.9 52.2 48.5 70.0 60.9 53.5 15.8 24.2 94.1
5   D2-Pruning 73.0 58.4 41.9 69.3 51.8 85.7 1391.2 65.7 57.6 63.9 62.1 52.5 15.3 22.3 94.8
6   COINCIDE 76.5 59.8 46.8 69.2 55.6 86.1 1495.6 63.1 54.5 67.3 62.3 53.3 16.1 24.3 97.8
7   Random 75.7 58.9 44.3 68.5 55.3 84.7 1483.0 62.2 54.8 65.0 61.7 50.2 15.1 21.9 95.0
8   CLIP-Score 73.4 51.4 43.0 65.0 54.7 85.3 1331.6 55.2 52.0 66.2 61.0 49.1 14.3 20.3 90.6
9   Perplexity 75.8 57.0 47.8 65.1 52.8 82.6 1341.4 52.0 45.8 68.3 60.8 48.7 14.5 20.9 91.1
PROGRESS
10   Loss as Obj. 75.7 58.6 49.6 70.1 55.1 86.3 1498.4 62.5 55.5 65.5 63.4 53.3 17.3 23.7 98.4
11    Accuracy as Obj. 75.2 58.8 53.4 69.9 55.1 85.9 1483.2 61.1 54.4 65.5 63.0 52.8 17.3 24.6 98.8
LLaVA-v1.5-13B
12   Full-Finetune 80.0 63.3 58.9 71.2 60.2 86.7 1541.7 68.5 61.5 69.5 68.3 60.1 19.3 22.1 100
13   Self-Sup 76.3 60.5 50.0 70.2 52.7 85.4 1463.8 63.7 57.6 64.9 65.2 53.3 17.2 23.2 93.8
14   Self-Filter 75.0 59.8 48.6 69.5 55.8 84.5 1446.9 58.8 51.8 69.1 65.3 52.4 16.9 23.1 92.6
15   EL2N 77.2 59.6 54.8 69.9 56.1 84.1 1531.0 59.3 52.3 65.8 65.7 53.9 17.0 24.4 94.4
16   SemDeDup 75.6 57.5 48.3 70.5 57.7 85.3 1397.6 59.0 51.1 68.7 64.9 53.2 16.8 24.6 92.9
17   D2-Pruning 73.9 60.5 49.8 70.4 55.2 84.9 1463.0 67.3 59.9 66.5 65.9 53.4 16.9 23.5 94.7
18   COINCIDE 77.3 59.6 49.6 69.2 58.0 87.1 1533.5 64.5 56.6 66.4 65.9 52.9 18.4 25.0 95.9
19   Random 76.7 60.5 48.0 68.8 57.7 84.8 1484.9 62.8 55.2 68.6 65.5 57.9 17.1 24.3 95.0
20   CLIP-Score 75.3 52.6 42.2 69.7 57.3 85.4 1426.3 60.4 54.0 68.1 63.3 52.8 17.4 23.7 91.8
21   Perplexity 77.0 58.5 48.2 68.7 54.8 83.1 1508.8 57.5 50.3 68.7 64.7 53.1 17.6 23.8 92.7
PROGRESS
22   Loss as Obj. 76.8 59.7 54.6 70.4 58.0 87.2 1458.3 63.8 56.9 69.9 65.1 58.0 17.9 24.6 96.8
23   Accuracy as Obj. 76.9 58.9 53.0 70.1 57.5 87.1 1497.6 63.9 57.6 67.3 65.4 57.7 18.0 24.5 96.5
@@ 4.2 Results and Analysis
PROGRESS is more effective than existing SOTA in data efficient learning. Table 1 (Row 0-11) compares PROGRESS against state-of-the-art baselines for training LLaVA-v1.5-7B on LLaVA-665K dataset under a 20% data budget, following standard settings. [^3] Reproduced with official code. PROGRESS achieves the highest relative performance (98.8%), outperforming all baselines, including those requiring access to ground-truth answers for the entire dataset and additional reference VLMs. In contrast, PROGRESS uses supervision only on a need basis for 20% of samples and relies solely on self-supervised features, yet reaching near-parity with full finetuning. PROGRESS also ranks among the top two methods on 8 out of 14 benchmarks, showing strong generalization across diverse tasks—e.g., including perception-focussed VQAv2 (75.2), scientific questions and diagrams (ChartQA:17.3, AI2D:52.8), and object hallucination POPE (85.9). Notably, it exceeds full-data performance on VizWiz (53.4 vs. 47.8), SQA-I (69.9 vs. 68.4), MME (1483.2 vs. 1476.9), ChartQA (17.3 vs. 16.4) and CMMMU (24.6 vs. 22.1). These results demonstrate effectiveness the PROGRESS as a dynamic and fully automated alternative for efficient VLM training under limited supervision.
@@
Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA- SEED Rel. (%)
en cn Wild
Architecture Generalization (Qwen2-VL-7B)
Full-Finetune 77.4 61.7 45.5 81.4 59.7 84.3 1567.9 76.1 75.1 84.8 66.9 100
Random 76.2 60.1 43.6 81.4 58.7 83.7 1556.8 76.8 74.5 81.7 67.6 98.7
COINCIDE 76.7 60.2 45.4 81.7 59.4 83.6 1583.5 77.4 76.2 80.5 67.9 99.6
PROGRESS 76.2 60.5 47.1 82.3 58.0 84.3 1560.1 77.2 72.9 87.1 67.6 100.0
Dataset Generalization (Vision-Flan-191K)
Full-Finetune 69.4 46.0 49.7 59.9 34.1 85.1 1306.1 49.1 51.7 35.7 53.3 100
Random 66.0 43.8 52.2 62.1 39.7 82.7 1072.2 48.7 43.7 40.4 28.7 95.0
COINCIDE 66.3 43.6 51.0 63.8 35.2 81.9 1222.2 56.7 45.5 31.1 37.5 95.8
PROGRESS 65.5 44.0 53.6 62.5 42.0 82.9 1040.9 43.6 47.4 43.2 45.3 99.0
@@
Architectures and Dataset Generalization. In Table 2, we test generalization of PROGRESS across different VLM architecture and IT dataset with accuracy as signal. For architecture generalization, we use newer Qwen2-VL-7B and train it on the LLaVA-665K dataset using the same 20% data budget and identical hyperparameters. We compare PROGRESS with two of the strongest (highest performing) established baselines—Random Sampling and COINCIDE—across multiple multimodal benchmarks. PROGRESS achieves the highest overall relative performance of 100% and ranks first or second on 9 out of 11 benchmarks (Tab. 2, top). For dataset generalization, we report LLaVA-v1.5-7B on Vision-Flan dataset under a stricter 16.7% annotation budget to assess generalization in low-resource settings. PROGRESS achieves the highest overall relative performance of 99.0%, outperforming COINCIDE (95.8%) and Random (95.0%) and ranks first or second on 8 out of 11 benchmarks (Tab. 2, bottom).
The scalability and generalization are achieved without requiring any model-specific tuning. These results underscore the calability and generalization of PROGRESS , making it a practical solution for efficient training across diverse architecture and datasets.
4.3 Investigating the effectivenessof different components of
PROGRESS
In this section, we provide further insights into the learning behaviour of PROGRESS . All experiments use LLaVA-v1.5-7B on the LLaVA-665K dataset with 20% sampling ratio and accuracy as the objective unless otherwise specified.
How important is enforcing Diversity among the selected samples? To assess the importance of balancing informativeness and diversity (Eqn 2), we conduct an ablation study varying the temperature parameter $\tau$ in the softmax used for skill selection (Eqn 2). As shown in Figure 4, a high temperature of 1.0 yields the best overall performance (98.8% relative score), striking an effective balance between prioritizing high-improvement clusters and maintaining diversity across concepts. As the temperature decreases (i.e., $\tau=0.7,0.5,0.3$), performance consistently degrades, with the lowest temperature yielding only 92.8%—a significant drop of over 6% in relative score. This decline confirms that overly sharp sampling distributions (low $\tau$) lead to mode collapse, where the model repeatedly focuses on a narrow set of concepts and fails to generalize broadly. Thus, we see that enforcing diversity helps improve performance. However, excessive diversity ($\tau$ = 1.2) is also not good as, in that case, the high-improvement clusters start losing their clear priority over other clusters.
Table 3: Ablation of Selection Policy. Performance comparison of different selection policies with the same warm-up strategy.
Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA- SEED AI2D ChartQA CMMMU Rel. (%)
en cn Wild
0   Full-Finetune 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 67.0 56.4 16.4 22.1 100
1   Warm-up Only 73.1 55.9 43.8 67.9 54.2 85.4 1410.3 58.5 52.7 64.6 60.5 52.4 16.1 24.5 94.6
2   Random 75.7 59.0 43.8 68.8 54.9 85.6 1414.2 61.9 54.9 66.2 63.3 48.6 17.3 25.2 96.8
3   Easiest 72.0 54.8 50.2 67.1 51.6 85.7 1407.4 57.0 52.6 65.2 59.5 50.1 12.3 22.8 92.3
4   Medium 69.3 52.5 46.0 68.3 50.8 85.4 1307.6 54.6 48.7 62.5 57.7 47.6 14.3 26.1 91.1
5   Hardest 72.8 54.8 52.1 61.3 50.5 85.4 1364.8 37.9 34.5 67.5 54.1 41.4 15.8 25.9 88.5
PROGRESS
6   Loss as Obj. 75.7 58.6 49.6 70.1 55.1 86.3 1498.4 62.5 55.5 65.5 63.4 53.3 17.3 23.7 98.4
7    Accuracy as Obj. 75.2 58.8 53.4 69.9 55.1 85.9 1483.2 61.1 54.4 65.5 63.0 52.8 17.3 24.6 98.8
@@ 4.3 Investigating the effectivenessof different components of
How effective is our Selection Policy? We evaluate the efficacy of PROGRESS (sampling based on relative accuracy change - Sec 3.2) by comparing it against other competitive selection strategies: Random Sampling, Easiest-Sampling (selecting clusters with highest absolute accuracy at given time step), Medium-Sampling (selecting moderate accuracy clusters), and Hardest-Sampling (selecting lowest accuracy clusters). As shown in Tab. 3, PROGRESS achieves the highest relative score (98.8%), ranking first on 8 out of 14 benchmarks and second on 4 others. To further analyze learning dynamics across selection strategies,
Figure 6: Ablation of Sampling Ratio. Relative performance on Vision-Flan dataset under different sampling ratio.
we categorize skill clusters into three difficulty levels—easy, moderate, and hard—based on their accuracy at the start of training. We then track the avg. performance of 50 clusters from each difficulty level for different selection strategies. As shown in Figure 5, PROGRESS consistently achieves higher mean performance and lower variance across all difficulty levels. These findings indicate that PROGRESS effectively balances learning across varying task difficulties, outperforming all other baseline strategies in both performance and stability.
How effective is PROGRESS under different sampling ratios? We show relative performance on the Vision-FLAN dataset under different sampling ratios (even lower than 16.7 % considered in Tab. 2) in Figure 6. PROGRESS consistently outperforms strongest baselines - Random and COINCIDE across different sampling ratios, highlighting its effectiveness in prioritizing informative samples.
How important is the order of skill acquisition? Here we randomly shuffle PROGRESS-selected samples and perform training —thereby ablating importance of learning order. We find a 4.1% drop in relative performance, underscoring the importance of introducing appropriate skills at the right time. More details are in Appendix B.
Wall Clock time - More details in Appendix B.
4.4 Digging Deeper into the Learning Behaviour of the Model
@@
Figure 7: Accuracy improvements with (a) benchmark difficulty and (b) sample rarity in the training data. The largest performance gains occur in mid-range of difficulty and low-mid range of rarity.
What skills model prioritize and when? We analyze what skills are learned during the course of training (especially, in the beginning and towards the end). We use standard benchmark MME-defined (Liang et al., 2024) categories as it provides interpretable fine-grain abilities. Results are shown in Figure 8, panel (a) presents grounded perceptual skills - scene, position, and count, while panel (b) includes language and symbolic skills - OCR, text translation, and code reasoning. Each cluster is assigned a dominant ability (see Appendix B for the assignment algorithm). We track the number of samples selected per ability (shown as bars) and monitor corresponding accuracy trends throughout training on 20% samples. We observe that the model consistently performs better on certain abilities than others. For example, scene recognition outperforms position understanding, and OCR yields higher accuracy than both text translation and code reasoning. Interestingly, as shown in Fig. 8(a), the model initially prioritizes count (at iteration 541), but accuracy does not improve — likely because it is too early in training to benefit from learning this skill. The model temporarily shifts focus and improves other abilities (i.e scene) , then revisits count and position abilities at iteration 738, where accuracy now increases and later stabilizes. In Fig. 8(b), OCR initially shows a decline in accuracy, but the model gradually selects more OCR-relevant samples. At iteration 738, it selects the most OCR samples, leading to the largest accuracy gain. Toward the end of training, the model increasingly focuses on more difficult abilities like code reasoning and text translation, with sample selection and performance rising after iteration 790. Notably, PROGRESS outperforms full-data fine-tuning performance (denoted by dashed line) on these challenging skills, despite using only 20% of data.
Figure 8: What skills model prioritize and when? We track the number of samples selected per ability (shown as bars) and monitor corresponding accuracy trends throughout training on 20% samples. The dashed horizontal lines show full-data fine-tuning performance for comparison.

### v2

@@ 4.1 Experimental Setup
Datasets and Models. To demonstrate effectiveness and generalizability across different scales of instruction-tuning (IT) data, we follow standard protocol [19] & conduct experiments on two IT datasets: large-scale LLaVA-665K [25] containing $\sim 0.6$ Million samples & Vision-Flan-191K [47]. For target VLMs, we primarily use LLaVA-v1.5-7B [25], following prior work, and additionally report results on LLaVA-v1.5-13B [25] to test scalability, and Qwen2-VL-7B [43] and Qwen2.5-VL-32B-Instruct [2] to test generalization to newer architectures.
Implementation Details. Following standard protocol [19, 45], we start from a pretrained model and instruction-tune it on the datasets using LoRA [14] with the official hyperparameters from LLaVA-1.5 and Qwen2-VL. For the accuracy variant, cluster-wise accuracy is estimated with an LLM judge that compares VLM outputs to ground-truth answers, though this step is not required for our loss-based variant. We strictly follows standard setup, evaluation protocols and metrics from prior work to ensure consistency and fair comparison [19, 45]. Additional implementation details are provided in Appendix 6.2.
Baselines. Consistent with previous work  [19], we compare PROGRESS against strong baselines spanning five major categories: (1) scoring function methods (CLIP-Score, EL2N[36], Perplexity [31]); (2) deduplication-based selection (SemDeDup [1]); (3) graph-based methods (D2-Pruning [29]); (4) external curriculum driven (Self-Filter [6]); and (5) concept-diversity approaches (COINCIDE [19], Self-Sup [40]). Our baselines span both active learning (denoted by AL) and curriculum learning methods (denoted by CL) for comprehensive comparison. Following previous work, we include Random—a competitive baseline shown to perform well due to its diversity—and Full-Finetune, representing the performance upper bound with full data/supervision training. Details for baselines follow previous work [19] (See Appendix 6.1 for further details). Evaluation Benchmark. Following prior work, we evaluate our approach on a diverse suite of 14 benchmarks with wide coverage of skills: perceptual reasoning (VQAv2 [11], VizWiz [12]), textual reasoning (TextVQA [39]), compositional reasoning (GQA [15]), object hallucinations (POPE [21]), multilingual understanding (MMBench-cn [26], CMMMU [10]), instruction-following (LLaVA-Bench[25]), fine-grained skills (MME [22], MMBench-en [26], SEED [20]), and scientific questions and diagrams (SQA-I [28], AI2D [16], ChartQA [32]).
@@
Evaluation Metrics. Following all prior work, we use standard evaluation metrics to ensure consistency and fair comparison. Specifically, we report average relative performance [19] across benchmarks to provide a unified measure of generalization. For each benchmark, relative performance is defined as: $\text{Rel.}=\left(\frac{\text{Model Performance}}{\text{Full Data Finetuned Performance}}\right)\times 100\%$. This normalizes differences in performance scale and difficulty across benchmarks, following prior work.
@@ 4.2 Results and Analysis
PROGRESS is more effective than existing SOTA in data efficient learning. Table 1 (Row 0-11) compares PROGRESS against state-of-the-art baselines for training LLaVA-v1.5-7B on LLaVA-665K dataset under a 20% data budget, following standard protocol. PROGRESS achieves the highest relative performance (98.8%), outperforming all baselines, including those requiring access to ground-truth answers for the entire dataset and additional reference VLMs. In contrast, PROGRESS uses supervision only on a need basis for 20% of samples and relies solely on model’s own progress, yet reaching near-parity with full finetuning. Beyond aggregate gains, PROGRESS also ranks among the top two methods on 8 out of 14 benchmarks, showing strong generalization across diverse tasks (Table 1 (Row 0-11))—e.g., including perception-focussed VQAv2 (75.2), scientific questions and diagrams (ChartQA:17.3, AI2D:52.8), and object hallucination POPE (85.9). Notably, it exceeds full-data performance on VizWiz (53.4 vs. 47.8), SQA-I (69.9 vs. 68.4), MME (1483.2 vs. 1476.9), ChartQA (17.3 vs. 16.4) and CMMMU (24.6 vs. 22.1). These results demonstrate PROGRESS is a dynamic and fully automated alternative for efficient VLM training.
@@
VLM Architectures and Dataset Generalization. In Table 2, we test generalization of PROGRESS across different VLM architecture and IT dataset with accuracy as signal. For architecture generalization, we use newer Qwen2-VL-7B and train it on the LLaVA-665K dataset using the same 20% data budget and identical hyperparameters. We compare PROGRESS with two of the strongest (highest performing) established baselines—Random Sampling and COINCIDE—across multiple multimodal benchmarks. PROGRESS achieves the highest overall relative performance of 100% and ranks first or second on 9 out of 11 benchmarks (Tab. 2, top). We also report results on Qwen2.5-VL-32B-Instruct (See Appendix 7.1), demonstrating the strong generalization to even larger-scale VLMs. For dataset generalization, we report LLaVA-v1.5-7B on Vision-Flan dataset under a stricter 16.7% annotation budget (standard protocol) to assess generalization in low-resource settings. PROGRESS achieves the highest overall relative performance of 99.0%, outperforming COINCIDE (95.8%) and Random (95.0%) and ranks first or second on 8 out of 11 benchmarks (Tab. 2, bottom). These results underscore the calability and generalization of PROGRESS , making it a practical solution for efficient training across diverse architecture and datasets.
4.3 Investigating the effectiveness of different components of PROGRESS
We analyze & ablate the components of PROGRESS . We use LLaVA-v1.5-7B on LLaVA-665K with 20% sampling and accuracy as the objective unless stated otherwise.
Table 3: Ablation of Selection Policy. Performance comparison of PROGRESS with baselines and curriculum learning (CL)-based selection strategies under identical warmup conditions.
[…]
How effective is our Selection Policy (i.e Eqn 1)? We evaluate our progress-based selection policy (Sec. 3.2) against several competitive strategies in Table 3. All strategies use identical warmup conditions. We compare with curriculum baselines: Random Selection, Easiest-Sampling (selecting clusters with highest absolute accuracy at a given step), Medium-Sampling (selecting mid-accuracy clusters), and Hardest-Sampling (selecting lowest-accuracy clusters). As shown in Tab. 3, PROGRESS achieves the highest relative score (98.8%), ranking first on 7 out of 14 benchmarks and second on 5 others. To further analyze learning dynamics across strategies, we group skill clusters into difficulty levels: easy, moderate, and hard based on their initial performance, and track the average performance of 50 clusters per level (see Fig. 4). PROGRESS consistently achieves higher mean performance with lower variance across all levels, effectively balancing learning across task difficulties throughout training. Row 2 includes a DINO–BERT version of COINCIDE, which replaces Tiny-LLaVA features with the same unsupervised features used in our method, confirming that the gains of PROGRESS arise from the progress-driven selection policy rather than the warmup or feature choice.
@@ 4.3 Investigating the effectiveness of different components of PROGRESS
How important is balancing informativeness & diversity in selected samples (i.e Eqn 2)? We ablate the temperature $\tau$ in the softmax skill-selection (Eqn. 2). As discussed in Sec 3.2, lower $\tau$ would over-focuses on top clusters, reducing diversity, while very high $\tau$ makes high-improvement clusters lose priority. Fig. 5 shows, $\tau=1.0$ achieves the best performance (98.8% Rel.), balancing priority and diversity. Decreasing $\tau$ (0.7, 0.5, 0.3) degrades performance, with lowest $\tau=0.3$ dropping to 92.8% (-6%), confirming that overly sharp distributions cause mode collapse. Excessive diversity ($\tau=1.2$) also hurts performance, as high-improvement clusters start to lose their clear priority during selection.
Ablation of Warmup. Table 4 shows that warmup phase alone provides only modest performance (94.6% Rel.), while PROGRESS yields large gains (+4% relative) over the warmup-only baseline (rows 2 vs. row 1). This demonstrates that warmup merely initializes reasonable per-cluster estimates, whereas the major improvements stem from PROGRESS ’s progress-driven sample selection. Row 3 shows PROGRESS remains robust to different warmup strategies, consistently delivering near-full-data performance even with weak (random sampling) warmup.
Table 4: Ablation of Warmup. Performance comparison of PROGRESS against warmup-only and across different warmup strategies. PROGRESS significantly boosts performance over warmup and remains robust across different warmup strategies.
@@
Time Efficiency and Annotation Cost Analysis. Consistent with prior work [19], we measure the wall-clock cost of the entire pipeline—data selection plus model finetuning—against relative performance (Rel.). As shown in Fig. 6, PROGRESS reaches relative performances of 96.3%, 97.1%, 98.2%, 99.1%, and 100% within 1.5, 1.75, 2.5, 4, and 5.67 hrs, respectively, making it
Figure 6: Wall-clock time comparison. We show results on LLaVA-665K here and Vision-Flan-191K in App. 7.3.
Pareto-superior to COINCIDE. Furthermore, Full-data finetuning requires $\sim$9 hr (and 100% data for training), much higher than our method which needs only 5.67hr of total training time and 20% data for training.
Note: Runtime includes all cost for entire pipeline, for fair comparison (Fig. 6), specifically: PROGRESS comprises feature extraction, K-means clustering, self-evaluation, warm-up and training, while COINCIDE includes feature extraction, clustering, and training. Both methods use same GPU compute and 20% data to train following standard protocol from prior work. See Appendix 7.4 for detailed stage-wise time breakdown. Beyond wall-clock gains, annotation cost dominates end-to-end cost in VLM training ($\approx$1902 hr for LLaVA-665K); by requiring only 20% labels, PROGRESS reduces annotation time by 80% ($\approx$380 hr), yielding the largest overall savings when combining annotation and training. See Appendix 7.5 details.
Figure 7: Ablation with different data budget for training. Relative performance on Vision-Flan dataset under different sampling ratio.
How effective is PROGRESS under different data budget? Fig 7 shows performance under different data budget used for training. PROGRESS consistently outperforms strongest baselines - Random and COINCIDE across different sampling ratios, highlighting its effectiveness. Notably, on scaling data size to higher percentages 32%, 64%, our method outperforms full-data-finetuning (which uses 100% data) by larger and larger margin. See Appendix 7.6 for more details on how scaling data improves PROGRESS more than Full data-Finetuning.
Representativeness of the Selected Subset. We assess how well each selected subset represents the full data distribution using PCA-based representativeness scores in the data feature space. PROGRESS yields a substantially lower representativeness distance (lower is better) to the full data pool (||0.083|| vs. ||0.198||), reflecting stronger global coverage. PROGRESS retains more variance than COINCIDE (38.7% vs. 34.2%), and captures a larger fraction of the top principal components (77–88% vs. 48–72%). These results show that PROGRESS spans the overall data distribution more broadly, while COINCIDE selects a narrower, more concentrated subset. See Appendix 7.2 for analysis details
Ablation Sensitivity to Hyperparameters:
[…]
Importance of skill order? See Appendix 7.2.
@@
Figure 8: Accuracy improvements with (a) benchmark difficulty and (b) sample rarity. Largest performance gains occur in mid-range of difficulty & low-mid range of rarity.
4.4 Analyzing Model Learning Behavior
How does benchmark difficulty and data frequency impact performance? In Fig. 8(a), we plot accuracy gains/improvement of PROGRESS as a function of benchmark difficulty, defined as the gap from full-data finetuned performance ((100full-finetune score)/100(100-$\text${full-finetune score})/100) (details in Appendix 7.7.1). PROGRESS achieves the largest gains on benchmarks of moderate difficulty. At the extremes, easy tasks such as POPE show limited improvements due to performance saturation, while hard tasks such as ChartQA, CMMMU yield smaller absolute gains because chart-related skills & rare Chinese-language skills are underrepresented in LLaVA-665K training dataset($\sim$0.96% & $\sim$1.1% respectively). Despite this scarcity, PROGRESS still surpasses full-data finetuning performance on both ChartQA and CMMMU with only 20% training data (Table 1, rows 10–11), showing that it can enhance performance of rare or niche abilities under limited supervision, though scarcity naturally constrains absolute gains. Fig. 8(b) further confirms this trend by plotting improvements against rarity, measured as $\log(1/\text{frequency})$ of benchmark-aligned samples (details in Appendix 7.7.2): once again, PROGRESS performs best in the mid-rarity regime, prioritizing skills that are neither over-abundant (with limited additional benefit from selecting more samples) nor rare (too few samples to generalize). Our findings align with the Zone of Proximal Development [42] & findings in [17], which state learning is most effective just beyond current ability where learning potential is highest—neither too easy nor too difficult.
(a) Language Abilities
(b) Perceptual Abilities
Figure 9: What skills model prioritize and when? We track the number of samples selected per ability (bars) and corresponding accuracy trends during training, with dashed lines indicating full-data finetuning performance.
What skills are prioritized during training? We analyze how different abilities emerge during training using fine-grained skill categories from prior work [22]. Fig. 9 groups clusters into language/symbolic (OCR, translation, code) abilities and grounded perceptual (scene, position, count). Each cluster is assigned a dominant ability (Appendix 7.8), and we track both selected samples and accuracy over time.
For language abilities, OCR, a prerequisite for translation and code reasoning, is strengthened first: an increase in OCR-selected samples in mid-training (iterations 738–790) coincides with OCR recovering from an early dip and exceeding full-data performance. Subsequent budget allocation to translation and code leads to steady accuracy gains that remain above full-data baselines. For perceptual abilities, scene—requiring coarse global recognition—acts as an easier prerequisite compared to count, which also demands precise localization and enumeration. The model initially prioritizes count (at iteration 541), but accuracy does not improve — likely because it is too early in training to benefit from learning this skill. It temporarily shifts focus and improves other abilities (i.e scene), then revisits count and position abilities at iteration 738, where accuracy now increases and later stabilizes. Overall, this shows PROGRESS consolidates easier, prerequisite abilities first and progressively invests its limited budget into harder skills.